Homogeneous differential equation


1. (xy^2)dx-(x^3+y^3)dy=0


2. (x^2+y^2)dx+xydy=0


3. (y^2-x^2)dx+2xydy=0


4. (3x+2y)dx-2xdy=0

1
Expert's answer
2021-11-22T15:09:19-0500

1.


"(xy^2)dy-(x^3+y^3)dx=0"

"(xy^2)y'-(x^3+y^3)=0"

"y'-\dfrac{1}{x}y=x^2y^{-2}"

First order Bernoulli ODE


"u=y^{1-(-2)}=y^3"

"u'=3y^2y'"

"\dfrac{1}{3}u'-\dfrac{1}{x}u=x^2"

"u'-\dfrac{3}{x}u=3x^2"

Integration factor


"\mu(x)=e^{\int(-3/x)dx}=x^{-3}"

"x^{-3}u'-\dfrac{3}{x^4}u=\dfrac{3}{x}""d(x^{-3}u)=\dfrac{3}{x}dx"

Integrate


"\int d(x^{-3}u)=\int\dfrac{3}{x}dx"

"x^{-3}u=\ln x+C"

"y^3=3x^3\ln x+Cx^3"

"y=x\sqrt[3]{3\ln x+C}"

2.


"(x^2+y^2)dx+xydy=0"

"x^2+y^2+xyy'=0"

"y'+\dfrac{1}{x}y=-xy^{-1}"

First order Bernoulli ODE


"u=y^{1-(-1)}=y^2"

"u'=2yy'"

"\dfrac{1}{2}u'+\dfrac{1}{x}u=-x"

"u'+\dfrac{2}{x}u=-2x"

Integration factor


"\mu(x)=e^{\int(2/x)dx}=x^2"

"x^2u'+2xu=-2x^3"

"d(x^2u)=-2x^3dx"

Integrate


"\int d(x^2u)=-\int2x^3dx"

"x^2u=-\dfrac{1}{2}x^4+C"

"y^2=-\dfrac{1}{2}x^2+\dfrac{C}{x^2}"

3.


"(y^2-x^2)dx+2xydy=0"

"y^2-x^2+2xyy'=0"

"y'+\dfrac{1}{2x}y=\dfrac{1}{2}xy^{-1}"

First order Bernoulli ODE


"u=y^{1-(-1)}=y^2"

"u'=2yy'"

"\dfrac{1}{2}u'+\dfrac{1}{2x}u=\dfrac{1}{2}x"

"u'+\dfrac{1}{x}u=xy^{-1}"

Integration factor


"\mu(x)=e^{\int(1/x)dx}=x"

"xu'+u=x^2"

"d(xu)=x^2dx"

Integrate


"\int d(xu)=\int x^2dx"

"xu=\dfrac{1}{3}x^3+C"

"y^2=\dfrac{1}{3}x^2+\dfrac{C}{x}"


4.


"(3x+2y)dx-2xdy=0"

"y'-\dfrac{1}{x}y=\dfrac{3}{2}"

Integration factor


"\mu(x)=e^{-\int(1/x)dx}=\dfrac{1}{x}"

"\dfrac{1}{x}y'-\dfrac{1}{x^2}y=\dfrac{3}{2x}"

"d(\dfrac{1}{x}y)=\dfrac{3}{2x}dx"

Integrate


"\int d(\dfrac{1}{x}y)=\int\dfrac{3}{2x}dx"

"\dfrac{1}{x}y=\dfrac{3}{2}\ln x+C"


"y=\dfrac{3}{2}x\ln x+Cx"